# Smooth Zone Barrier Lyapunov Functions for Nonlinear Constrained Control Systems

### v1

@@ Abstract
This paper introduces the Smooth Zone Barrier Lyapunov Function (s-ZBLF) for output and full-state constrained nonlinear control systems. Unlike traditional BLF methods, where control effort continuously increases as the state moves toward the constraint boundaries, the s-ZBLF method keeps the control effort nearly zero near the origin, with a more aggressive increase as the system approaches the boundary. However, unlike previous works where control effort was zero within a predefined safe region around the origin, the s-ZBLF overcomes the disadvantage of discontinuous control activation by providing a smooth, gradual increase in control effort as the state nears the constraints. This smooth transition improves continuity in the control response and enhances stability by reducing chattering. Additionally, the s-ZBLF provides the advantage of minimal control effort in regions far from the constraints, reducing energy consumption and actuator wear. Two forms of the s-ZBLF—logarithmic-based and rational-based—are presented. Theoretical analysis guarantees that all system states remain within the defined constraints, ensuring boundedness and stability of the closed-loop system. Simulation results validate the effectiveness of the proposed method in handling constrained nonlinear systems.
@@
Zone Barrier Lyapunov Function, Constrained Control, Nonlinear System, Adaptive Control, Full State Constraints.
@@ I Introduction
To overcome these limitations, the Zone Barrier Lyapunov Function (zBLF) was introduced, offering an improvement over traditional BLFs by introducing the concept of zones [4]. In zBLF, system states are allowed to move freely within predefined safe zones, and control effort is only applied as the states approach the boundary of these zones. This significantly reduces unnecessary control actions, improves energy efficiency, and reduces actuator wear. The zBLF method has shown great promise, particularly in applications where maintaining efficient control is important, as it allows the system to operate without intervention within safe zones.
Despite the advantages of zBLF, there are some inherent limitations. Most notably, zBLF suffers from discontinuous control activation near the boundary. As the system states transition from the free-moving zone to the constrained region, the control effort can switch on abruptly, leading to sharp control actions. This lack of fine-grained control near the boundaries can result in chattering, control instability, or degraded system performance, particularly in applications requiring smooth and continuous control. Additionally, zBLF’s reliance on strict boundaries can make it challenging to ensure a smooth and gradual increase in control effort as the system approaches the constraints.
To address these challenges, we propose a new approach called the Smooth Zone Barrier Lyapunov Function (s-ZBLF). The s-ZBLF retains the energy-saving advantages of zBLF while addressing the issue of discontinuous control activation. Our approach introduces a smooth transition between the safe zone and the constrained region, providing a gradual increase in control effort as the system state moves closer to the boundary. By making the control effort grow progressively, s-ZBLF ensures continuity in the Lyapunov derivative, which leads to smoother and more stable control behavior. Near the origin, where the error is low, the control effort remains minimal, gradually increasing as the system approaches the constraint boundaries, and applying hard constraints only when the boundary is near. This progression ensures that control actions are smoother, leading to more effective constraint enforcement while maintaining system stability.
@@
  • We propose the s-ZBLF, which introduces smooth transitions in control effort as the system approaches constraint boundaries, addressing the discontinuous control issue present in zBLF.
  • s-ZBLF maintains minimal control effort when the system state is far from the constraint boundaries, while gradually increasing the control effort as the state moves closer to the constraints, ensuring smooth and stable transitions.
  • The proposed method is demonstrated through theoretical analysis and simulation results, showing the performance of smooth control, stability, and constraint handling of s-ZBLF.
@@ II Problem Formulation and Preliminaries
$\dot{x}_{i}=f_{i}(x_{1},\dots,x_{i})+g_{i}(x_{1},\dots,x_{i})x_{i+1},\quad i=1,\dots,n-1,$ (1)
$\dot{x}_{n}=f_{n}(x_{1},\dots,x_{n})+g_{n}(x_{1},\dots,x_{n})u,$ (2)
@@
Control Objective: The goal is to design a control law $u(t)$ such that the system tracks the desired trajectory $x_{d}(t)$, while ensuring the state constraints $|x_{i}(t)|<k_{xi}$ are satisfied, with $i=1$ for output constraints and $i=1,\dots,n$ for full-state constraints. Additionally, all closed-loop signals must remain bounded for $t\geq 0$.
@@
There are several types of BLF, such as tangent-type [5], integral-type [6], and logarithmic-type [7]. One example of a standard BLF is given by the logarithmic form $V(x)=\frac{1}{2}\ln\left(\frac{k^{2}}{k^{2}-x^{2}}\right)$, where $k$ represents the constraint boundary.
@@
III Control Design with Smooth Zone Barrier Lyapunov Function
III-A Smooth Zone Barrier Lyapunov Function
Smooth Zone Barrier Lyapunov Function is introduced to provide smooth and continuous control activation, ensuring minimal control effort near the origin and progressively increasing control effort as the state nears the boundary. Unlike ZBLF, where control activation may be abrupt, s-ZBLF provides a gradual increase in control effort as the system state approaches the boundary, eliminating discontinuities and enhancing system stability. For a system state $x$ with constraint $|x|<k$, the following two examples of s-ZBLF are considered:
@@ III-B Motivation Example: Control Design for a Second Order SISO System
where $u$ is the control input. Let $z_{1}=x_{1}-x_{d}$ and $z_{2}=x_{2}-\alpha_{1}$, where $\alpha_{1}$ is a stabilizing function. The objective is to design an output constrained control law $u$ that respects the state constraint $|x_{1}(t)|<k_{x1}$ while maintaining system stability. We illustrate both forms of the s-ZBLF presented in the previous section with examples.
III-B1 Logarithmic-Based s-ZBLF
@@
III-B2 Rational-Based s-ZBLF
For the rational-based s-ZBLF, the Lyapunov function is defined as:
@@ III-B2 Rational-Based s-ZBLF
This ensures $\dot{V}_{2}\leq 0$, establishing the stability of the closed-loop system, and the boundedness of $x_{1}$ considering Assumption 1, and Lemma 2 for the rational-based s-ZBLF.
III-C s-ZBLF for Higher-Order Nonlinear Systems with Output Constraints
Backstepping is used in the control design in this section. Let $z_{1}=x_{1}-x_{d}$ and $z_{i}=x_{i}-\alpha_{i-1}$, for $i=2,\dots,n$. Define the state constraint as $|x_{1}(t)|<k_{x1}$, and let $k_{1}=k_{x1}-k_{xd}$ is the error constraint. The Lyapunov function candidates are defined using the logarithmic-based s-ZBLF as follows:
@@
III-D s-ZBLF for Higher-Order Nonlinear Systems with Full State Constraints
@@ IV Simulation
where $\Theta_{1}=0.1$, $\Theta_{2}=0.1$, and $\Theta_{3}=-0.2$ [1]. The control design employs a s-ZBLF for output constraints, with $k_{1}=0.56$. The objective is to track the desired trajectory $y_{d}=0.2+0.3\sin(t)$, ensuring that the state $x_{1}$ remains within the safe constraint $|x_{1}|<k_{1}$. The initial conditions are set as $x_{1}(0)=0.25$ and $x_{2}(0)=1.5$. The control parameters $\kappa_{1}=2$ and $\kappa_{2}=2$ are used to adjust the response, while $\beta=10$ governs the control effort transition near the boundaries.
@@ V Conclusion
In this paper, we introduced the s-ZBLF for output and full-state constrained nonlinear control systems. The method offers a novel approach by maintaining minimal control effort near the origin and smoothly increasing it as the system state moves toward the constraint boundaries, providing a continuous and stable control response. This gradual transition avoids abrupt changes in control effort, addressing limitations found in earlier methods, and improves system performance by reducing chattering. Two specific forms—logarithmic-based and rational-based—were proposed, each demonstrating their effectiveness through theoretical analysis and simulation results for output-constrained systems. Future work will focus on optimizing the parameter $\beta$ to fine-tune the rate of control effort increase as the system state approaches the constraints. Additionally, we will explore the extension of s-ZBLF to multi-input multi-output (MIMO) systems and adaptive control schemes, as well as its application to systems with partial, asymmetric, time-varying, or unknown constraints to further enhance its practical applicability.

### v2

# Progressive Barrier Lyapunov Functions for Nonlinear Constrained Control Systems

@@ Abstract
This paper introduces the Progressive Barrier Lyapunov Function (p-BLF) for output- and full-state-constrained nonlinear control systems. Unlike traditional BLF methods, where control effort continuously increases as the state approaches the constraint boundaries, the p-BLF maintains minimal control effort in unconstrained regions and increases it progressively toward the boundaries. In contrast to previous methods with predefined safe zones and abrupt control activation, the p-BLF provides a smooth transition, improving continuity in the control response and enhancing stability by reducing chattering. Two forms of the p-BLF, logarithmic-based and rational-based, are developed to handle systems with either output constraints or full-state constraints. Theoretical analysis guarantees that all system states remain within the defined constraints, ensuring boundedness and stability of the closed-loop system. Simulation results validate the effectiveness of the proposed method in constrained nonlinear control.
@@
Progressive Barrier Lyapunov Function, Constrained Control, Nonlinear System, Adaptive Control, Full State Constraints.
@@ I Introduction
Traditional BLF-based control strategies have been extensively studied for enforcing state constraints in nonlinear systems. Various forms of BLFs have been proposed, including tangent-type [9], integral-type [5], and logarithmic-type functions [1, 6], all designed to increase rapidly as the state approaches the constraint boundary. While this property ensures strong constraint enforcement, it also leads to high control effort even when the system states remain far from the boundaries. Such unnecessary control activation can increase energy consumption and accelerate actuator wear, particularly in applications where the constraints are rarely active.
To overcome the high control effort associated with traditional BLFs, the Zone Barrier Lyapunov Function (zBLF) was introduced as an alternative that defines a predefined safe region around the origin [4, 8]. Within this zone, system states are allowed to evolve freely without control activation, and corrective effort is applied only as the state approaches the zone boundary. This structure helps reduce unnecessary control actions, improve energy efficiency, and limit actuator usage. However, zBLF presents limitations related to its discontinuous control behavior at the zone boundary. The abrupt activation of control introduces nonsmoothness in the system’s response, which can lead to discontinuities in the velocity and sharp variations in acceleration. These effects may cause chattering, degraded tracking performance, and increased mechanical stress in applications that require smooth and continuous dynamics. Moreover, the fixed zone definition complicates the design of controllers that aim to provide a gradual and continuous increase in effort as the constraint boundary is approached.
To address these challenges, we propose a new class of BLFs, referred to as the Progressive Barrier Lyapunov Function (p-BLF). The p-BLF method retains the low-effort behavior of zone-based approaches but removes the need for explicit zone definitions. It provides a gradual increase in control effort as the system state approaches the constraint boundaries, avoiding nonsmooth control activation. This progression ensures continuity in the Lyapunov derivative and improves stability by reducing chattering and discontinuities in the system response. The control effort remains minimal in unconstrained regions and increases only as needed to enforce the constraints.
@@
  • A new class of BLFs, called p-BLF, is introduced to enable gradual and continuous control activation near constraint boundaries, addressing nonsmoothness and discontinuities in existing methods.
  • The p-BLF framework maintains minimal control effort in unconstrained regions and increases it progressively as the state approaches the constraints, improving control efficiency and smoothness.
  • Theoretical analysis and simulations demonstrate that the proposed approach ensures constraint satisfaction, closed-loop stability, and continuous control performance.
The rest of the paper is structured as follows. Section II presents the problem formulation and relevant preliminaries. Section III introduces the proposed p-BLF, detailing its logarithmic-based and rational-based forms, and presents a control design example for a second-order SISO system. The section further extends the control design to higher-order systems with output and full-state constraints. Section IV presents simulation results, and Section V concludes the paper.
@@ II Problem Formulation and Preliminaries
$\dot{x}_{i}$ $=f_{i}(x_{1},\dots,x_{i})+g_{i}(x_{1},\dots,x_{i})x_{i+1},$ $i=1,\dots,n-1,$
$\dot{x}_{n}$ $=f_{n}(x_{1},\dots,x_{n})+g_{n}(x_{1},\dots,x_{n})u,$
$y$ $=x_{1}.$ (1)
@@
Control Objective: Given the nonlinear system described by (II), the objective is to design a control law $u(t)$ such that the output $y(t)=x_{1}(t)$ tracks a bounded desired trajectory $x_{d}(t)$, while satisfying the state constraints. Two scenarios are considered: (i) output constraint $|x_{1}(t)|<k_{x1}$, and (ii) full-state constraints $|x_{i}(t)|<k_{xi}$, $i=1,\dots,n$. To achieve this, a p-BLF is developed and integrated into the control design to ensure constraint satisfaction, closed-loop stability, and asymptotic tracking, while maintaining boundedness of all signals for $t\geq 0$.
@@
One example of a standard BLF is given by the logarithmic form $V(x)=\frac{1}{2}\ln\left(\frac{k^{2}}{k^{2}-x^{2}}\right)$, where $k$ represents the constraint boundary.
@@
III Control Design with Progressive Barrier Lyapunov Function
III-A Progressive Barrier Lyapunov Function
The Progressive Barrier Lyapunov Function (p-BLF) is introduced to enable smooth and continuous control effort modulation. It ensures minimal control activation near the origin and progressively increases the effort as the system state approaches the constraint boundary. In contrast to zone-based BLFs, which may cause abrupt control switching, the p-BLF provides a gradual transition, improving continuity and closed-loop stability.
For a system state $x$ subject to the constraint $|x|<k$, two example forms of the p-BLF are considered:
@@ III-B Motivation Example: Control Design for a Second Order SISO System
where $u$ is the control input. Let $z_{1}=x_{1}-x_{d}$ and $z_{2}=x_{2}-\alpha_{1}$, where $\alpha_{1}$ is a stabilizing function. The objective is to design an output constrained control law $u$ that respects the state constraint $|x_{1}(t)|<k_{x1}$ while maintaining system stability. We illustrate both forms of the p-BLF presented in the previous section with examples.
III-B1 Logarithmic-Based p-BLF
@@
III-B2 Rational-Based p-BLF
For the rational-based p-BLF, the Lyapunov function is defined as:
@@ III-B2 Rational-Based p-BLF
This ensures $\dot{V}_{2}\leq 0$, establishing the stability of the closed-loop system, and the boundedness of $x_{1}$ considering Assumption 1, and Lemma 2 for the rational-based p-BLF.
III-C p-BLF for Higher-Order Nonlinear Systems with Output Constraints
Backstepping is used in the control design in this section. Let $z_{1}=x_{1}-x_{d}$ and $z_{i}=x_{i}-\alpha_{i-1}$, for $i=2,\dots,n$. Define the state constraint as $|x_{1}(t)|<k_{x1}$, and let $k_{1}=k_{x1}-k_{xd}$ is the error constraint. The Lyapunov function candidates are defined using the logarithmic-based p-BLF as follows:
@@
III-D p-BLF for Higher-Order Nonlinear Systems with Full State Constraints
@@ IV Simulation
where $\Theta_{1}=0.1$, $\Theta_{2}=0.1$, and $\Theta_{3}=-0.2$ [1]. The control design employs a p-BLF for output constraints, with $k_{1}=0.56$. The objective is to track the desired trajectory $y_{d}=0.2+0.3\sin(t)$, ensuring that the state $x_{1}$ remains within the safe constraint $|x_{1}|<k_{1}$. The initial conditions are set as $x_{1}(0)=0.25$ and $x_{2}(0)=1.5$. The control parameters $\kappa_{1}=2$ and $\kappa_{2}=2$ are used to adjust the response, while $\beta=10$ governs the control effort transition near the boundaries.
@@ V Conclusion
In this paper, we introduced the p-BLF for output and full-state constrained nonlinear control systems. The method offers a novel approach by maintaining minimal control effort near the origin and smoothly increasing it as the system state moves toward the constraint boundaries, providing a continuous and stable control response. This gradual transition avoids abrupt changes in control effort, addressing limitations found in earlier methods, and improves system performance by reducing chattering. Two specific forms—logarithmic-based and rational-based—were proposed, each demonstrating their effectiveness through theoretical analysis and simulation results for output-constrained systems. Future work will focus on optimizing the parameter $\beta$ to fine-tune the rate of control effort increase as the system state approaches the constraints. Additionally, we will explore the extension of p-BLF to multi-input multi-output systems and adaptive control schemes, as well as its application to systems with partial, asymmetric, time-varying, or unknown constraints to further enhance its practical applicability.